Ball A collide head on elastically with identical ball B (initially at rest) with speed 2 m /s then find the final velocity of Ball A and Ball B.
1
Expert's answer
2019-01-17T08:10:11-0500

Let velocity of Ball A directed in positive direction.

v_a=2 m/s,v_b=0 m/s.

From the conservation of momentum

mv_a=-mu_a+mu_b

u_b=v_a+u_a

From the conservation of energy

(mv_a^2)/2=(mu_a^2)/2+(mu_b^2)/2

v_a^2=u_a^2+(v_a+u_a )^2

The final velocity of Ball A

u_a=0 m/s.

The final velocity of Ball B

u_b=v_a+0=v_a=2 m/s.